OpenStudy (anonymous):

I need help figuring Identities.... cosx(tan^2x+1) = secx

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 1+tan^2(\theta)=sec^2(\theta)}}\quad thus \\ \quad \\ cos(x)[tan^2(x)+1]\implies ?\)

OpenStudy (anonymous):

thats where i'm having a problem with while i'm solving it....

OpenStudy (jdoe0001):

well... what would that give you anyway? keep in mind secant = 1/cosine

OpenStudy (anonymous):

i dont know about the cosine but i do know whats in the parenthesis. cosx(sec^2x)

OpenStudy (jdoe0001):

yeap thus

OpenStudy (phi):

I usually change everything to sin and cos and see if things simplify

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 1+tan^2(\theta)=sec^2(\theta)}}\quad thus \\ \quad \\ cos(x)[tan^2(x)+1]\implies cos(x)sec^2(x)\implies \cancel{ cos(x) }\cfrac{1}{\cancel{ cos^2(x) }}\) yeap... that'd work too

OpenStudy (anonymous):

thus . . . . i dont know what to do after that.

OpenStudy (jdoe0001):

http://www.mathwords.com/t/trig_identities.htm check what 1/cos(x) is then

OpenStudy (anonymous):

its secx. but how can i just get sex by itself. do i just cancel or something.

OpenStudy (anonymous):

secx*

OpenStudy (alyssa_xo):

That's what he did in the last step. \(\large cos(x)*sec^2(x)=\cancel{cos(x)}*\frac{1}{cos^\cancel{2}(x)}=\frac{1}{cosx}=secx\)

OpenStudy (alyssa_xo):

\(\large \frac{a^m}{a^n}=a^{m-n}\)

OpenStudy (anonymous):

Ah ok. Thanks